The tabel is posted twice. Hopefully it is clear enough in one of those photos.
Question6-10 are based on the following table Economy Prob Recession Below aV Average Above avg0 0.4 0.2 0.2 17% 3.0% 1,0% 3.0% 9.0% -6. 0% -1.0% 4,0% 8.0% 11.0% -1,0% 0.0% 1,0% 2.0% 8.0% Boom 0.1
7-What are the expected risk levels (standard deviations) of YY? a. 9.53% b, 2.60%) c. 6.02% d. None of the above 8-What are the expected coefficients of variation of the above equities (XX, ZZ, YY)? b. 3.25 c. 60.24 d. None of the above 9.What is the best equity to invest in? d. Data provided is not enough 10-What is the expected return of a portfolio that has 50% XX, 25% ZZ and 25% YY? a.-2.775% b. 6.325% C. 11.885% d. None of the above
Question6-10 are based on the f EconomyProb. Recession Below avg Average Above avg Boom 0.4 0.2 0.2 -17%) -3.0%) 1.0% 3.0% 9.0% -6.0% -1.0% 4.0%) 8.0 % 11.0% -1.0% 0.0%) 1.0%) 2.0%) 8.0%)
Answer #1
7…
Mean return for YY=Sum (prob.*exp.Return)
(0.4*-1%)+(0.2*0)+(0.2*1%)+(0.1*2%)+(0.1*8%)=
0.8%
Std. Devn. Of returns of YY=Sq. Rt. Of( Sum Prob. *(exp.ret.-Mean ret.)^2))
Sum Prob. *Variances =(0.4*(-1%-0.8%)^2)+(0.2*(0-0.8%)^2)+(0.2*(1%-0.8%)^2)+(0.1*(2%-0.8%)^2)+(0.1*(8%-0.8%)^2)=
0.000676
Std. devn.=Sq. Rt. Of variance=(0.000676)^(1/2)=
2.60%
ANSWER: b. 2.60%
8..
Mean return for XX=
(0.4*-17%)+(0.2*-3%)+(0.2*1%)+(0.1*3%)+(0.1*9%)=
-6.00%
Std. Devn. Of returns of YY=Sq. Rt. Of( Sum Prob. *Variances)
Sum Prob. *Variances =(0.4*(-17%-(-6%))^2)+(0.2*(-3%-(-6%))^2)+(0.2*(1%-(-6%))^2)+(0.1*(3%-(-6%))^2)+(0.1*(9%-(-6%))^2)=
0.00906
Std. devn.=(0.00906)^(1/2)=
9.52%
Mean return for ZZ=
(0.4*-6%)+(0.2*-1%)+(0.2*4%)+(0.1*8%)+(0.1*11%)=
0.10%
Std. Devn. Of returns of YY=Sq. Rt. Of( Sum Prob. *Variances)
Sum Prob. *Variances =(0.4*(-6%-0.1%)^2)+(0.2*(-1%-0.1%)^2)+(0.2*(4%-0.1%)^2)+(0.1*(8%-0.1%)^2)+(0.1*(11%-0.1%)^2)=
0.003629
Std. devn.=(0.003629)^(1/2)=
6.02%
Coefficient of Variation(oeff.COV)=Std. devn./Mean returns
Coeff.OV(XX)=9.52%/-6%
-1.59
Coeff.OV(YY)=2.60%/0.8%=
3.25
Coeff.OV(ZZ)=6.02%/0.1%=
60.2
ANSWER: -1.59, 60.2, & 3.25
9. Best equity to Invest in
c. YY
As the mean return is the highest &
also the std. devn. Of returns from the mean is the least
as shown in the foll.table
Summary XX YY ZZ
Mean -6.00% 0.8% 0.10%
Std. Devn. 9.52% 2.60% 6.02%
Coeff. OV -1.59 3.25 60.2
10.Expected return of the given portfolio=
Sum(Wt.*Mean Returns)
ie.(50%*-6%)+(25%*0.1%)+(25%*0.8%)=
-2.775%
ANSWER: a. -2.775%